# Generative modelling with jump-diffusions

### v1

@@ Appendix D The multivariate generalized Laplace distribution
I briefly review basic properties of the multivariate generalized Laplace distribution following [9].
@@
$G(\mathbf{k})=\left(\frac{1}{1+\frac{1}{2}\mathbf{k}^{\rm T}\boldsymbol{\Sigma}\mathbf{k}-i\boldsymbol{\mu}{\cdot}\mathbf{k}}\right)^{\nu},\qquad\qquad\mathbf{k}\in\mathbb{R}^{d}$ (68)
@@
[…]
@@
Since we are interested in describing the jump amplitudes in the noise process $\boldsymbol{\xi}(t)$, Eq. (6), we want no contribution to the drift leading to the constraint $\boldsymbol{\mu}=\mathbf{0}$. We also require an isotropic distribution, which is realized by setting $\boldsymbol{\Sigma}=\sigma^{2}\mathbf{I}$. The resulting distribution can be called the isotropic multivariate generalized Laplace distribution $\text{GL}(\sigma^{2},\nu)=\text{GAL}(\mathbf{0},\sigma^{2}\mathbf{I},\nu)$ and has the characteristic function
$G(k)=\left(1+\frac{\sigma^{2}}{2}k^{2}\right)^{-\nu}$ (71)
@@
Amplitude samples can be easily generated using Eq. (69): $\mathbf{L}\sim\text{GL}(\sigma^{2},\nu)$ has the representation
[…]
where $\mathbf{X}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ and $\Gamma\sim\text{Gamma}(\nu)$.
Appendix E Alternative derivation of the Liouville equation for $\tilde{p}$
@@
~$\tilde{\mathcal{L}}u(\mathbf{x})$ $=$ $-\mathbf{f}(\mathbf{x},T-t){\cdot}\nabla u(\mathbf{x})$ (73)
@@
~$\tilde{\mathcal{L}}u(\mathbf{x})$ $=$ $-\mathbf{f}(\mathbf{x},T-t){\cdot}\nabla u(\mathbf{x})+\lambda\,g(T-t)\int\mathrm{d}\mathbf{z}\,\rho(z)\int_{0}^{1}\mathrm{d}\gamma\,\left[\mathbf{z}{\cdot}\nabla u(\mathbf{x}+\gamma g(T-t)\mathbf{z})\right]\times$ (74)
@@
~$\tilde{\mathcal{L}}^{\dagger}\tilde{p}(\mathbf{x},t)$ $=$ $-\nabla{\cdot}\left(-\mathbf{f}(\mathbf{x},T-t)+\lambda\,g(T-t)\int_{0}^{1}\mathrm{d}\gamma\int\mathrm{d}\mathbf{z}\,\mathbf{z}\,\rho(z)\right.\times$ (76)
@@
Appendix F Numerical experiment
The implementation of the test is done in analogy to that in [14] so that values are comparable. The only difference is that the MSLE values for JDL-SDE and JDL-ODE are averaged over the two components, while the values from [14] are evaluated over the first component only. Due to the isotropy of $p_{\rm data}$ the two MSLE procedures lead to comparable values. 32,000 samples from $p_{\rm data}$ have been used for estimation of the generalized score function and 25,000 samples have been generated for testing. The functions $G_{1,2}(x,t)$ that determine the generalized score function via Eqs. (16,18) are evaluated by discretizing $t\in[0,10]$ and $x\in[0,10]$ into 200 equally spaced values and creating a function representation by interpolation. The neural network parametrizing $\mathbf{s}_{\boldsymbol{\theta}}$ has been kept simple, consisting of 4 layers with 200 elementwise $\tanh$ activation nodes. The weight in the loss Eq. (12) is taken as $w(t)=1/T$. The JDL implementation in Eqs. (21,23,24) uses $D=1$, $\lambda=1$, $\sigma^{2}=2$, $T=10$, $\Delta t=0.1$. For further details on the implementations of DLPM, LIM, DDPM, please see [14].

### v2

@@ Appendix D The multivariate Laplace distribution
I briefly review basic properties of the multivariate Laplace distribution following [11].
@@
$\chi(\mathbf{k})=\left(\frac{1}{1+\frac{1}{2}\mathbf{k}^{\rm T}\boldsymbol{\Sigma}\mathbf{k}-i\boldsymbol{\mu}{\cdot}\mathbf{k}}\right)^{\nu},\qquad\qquad\mathbf{k}\in\mathbb{R}^{d}$ (84)
@@
    $\mathbf{L}\overset{d}{=}\boldsymbol{\mu}\Gamma+\sqrt{\Gamma}\,\mathbf{X}$ (85)
@@
Since we are interested in describing the jump amplitudes in the noise process $\boldsymbol{\xi}(t)$, Eq. (6), we want no contribution to the drift leading to the constraint $\boldsymbol{\mu}=\mathbf{0}$. We also require an isotropic distribution, which is realized by setting $\boldsymbol{\Sigma}=\sigma^{2}\mathbf{I}$.
For the JL model only the $\nu=1$ value is relevant leading to the isotropic multivariate generalized Laplace distribution $\text{L}_{d}(\sigma^{2})=\text{GAL}(\mathbf{0},\sigma^{2}\mathbf{I},1)$, which has the characteristic function
$\chi_{\rm L}(\mathbf{k})=\left(1+\frac{\sigma^{2}}{2}k^{2}\right)^{-1}$ (87)
@@
Amplitude samples can be easily generated using Eq. (85): $\mathbf{L}\sim\text{L}_{d}(\sigma^{2})$ has the representation
$\mathbf{L}\overset{d}{=}\sqrt{\Gamma}\,\mathbf{X}$ (88)
where $\mathbf{X}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ and $\Gamma\sim\text{Gamma}(1)$.
Note that the PDFs associated with the characteristic function Eqs. (84,87) have closed analytical forms [11]. For Eq. (87) one obtains the isotropic jump PDF
$\rho_{\rm L}(\mathbf{z},\sigma)$ $=$ $\mathcal{F}^{-1}\left\{\left(1+\frac{\sigma^{2}}{2}k^{2}\right)^{-1}\right\}$ (89)
$=$ $\frac{2}{\sigma^{d}(2\pi)^{d/2}}\left(\frac{z}{\sigma\sqrt{2}}\right)^{1-\frac{d}{2}}K_{1-\frac{d}{2}}\left(\frac{z}{\sigma}\sqrt{2}\right)$
where $K_{\mu}$ denotes the modified Bessel function of the 2nd kind.
Appendix E Distribution of the jump increments $\tilde{\mathbf{J}}(t)$
From the definition Eq. (28) we can directly derive the characteristic function $\chi(\mathbf{k},t)$ of $\tilde{\mathbf{J}}(t)$ by evaluating the expected value:
$\chi(\mathbf{k},t)=\mathbb{E}\left[\exp\left\{i\mathbf{k}\sum_{j=1}^{N_{t}}\mathbf{A}_{j}e^{(t-\tau_{j})/2}\right\}\right]$ (90)
where $N_{t}\sim\text{Poisson}(\lambda t)$, $\tau_{j}\sim\text{U}([0,t])$, and $\mathbf{A}_{j}\sim\text{L}_{d}(\sigma^{2})$ i.i.d.. Calculating the expected value over $N_{t},\tau$ is straightforward
$\chi(\mathbf{k},t)$ $=$ $\exp\left\{\lambda\int_{0}^{t}\mathrm{d}s\left(\mathbb{E}\left[e^{i\mathbf{k}\mathbf{A}e^{(t-s)/2}}\right]-1\right)\right\}$ (91)
$=$ $\exp\left\{\lambda\int_{0}^{t}\mathrm{d}s\,\phi_{\rm L}\left(k\,e^{s/2}\right)\right\}$
using Eq. (8). Direct integration with Eq. (70) yields
$\int_{0}^{t}\mathrm{d}s\,\phi_{\rm L}\left(k\,e^{s/2}\right)=-\log\left(\frac{1+\frac{\sigma^{2}}{2}k^{2}e^{t}}{1+\frac{\sigma^{2}}{2}k^{2}}\right)$ (92)
The characteristic function is then
$\chi(\mathbf{k},t)$ $=$ $\frac{1+\frac{\sigma^{2}}{2}k^{2}}{1+\frac{\sigma^{2}}{2}k^{2}e^{t}}$ (93)
$=$ $\frac{1}{1+\frac{\sigma^{2}}{2}k^{2}e^{t}}(1-e^{-t})+e^{-t}$
As a consequence, the PDF of $\tilde{\mathbf{J}}(t)$ is given by
$p(\tilde{\mathbf{J}},t)=\tilde{\rho}_{\rm L}(\|\tilde{\mathbf{J}}\|,\sigma\,e^{t/2})(1-e^{-t})+\delta(\tilde{\mathbf{J}})e^{-t}$ (94)
and can be sampled as outlined in Sec. 3.3.
Appendix F Details on the numerical experiments
Two data distributions are considered: (a) a swiss roll distribution standardized to zero mean and unit variance; (b) a mixture of 9 equal Gaussians with variance $\sigma^{2}=0.05^{2}$ arranged in a grid with weights
$\mathbf{W}=\left(\begin{matrix}0.3&0.15&0.05\\ 0.1&0.02&0.15\\ 0.01&0.2&0.02\end{matrix}\right)$ (95)
The Gaussian mixture distribution (b) is adapted from [20].
Score functions are estimated by minimizing the DSM loss functions Eq. (25) (JL-ODE/SDE), Eq. (96) (Gauss-ODE), and Eq. (100) (LIM-SDE) using $10^{5}$ training samples. The neural network architecture is a simple multilayer perceptron with 4 layers of 124 nodes and GELU activation functions. Training is performed using an ADAM optimizer, a learning rate of $\approx 10^{-3}$, and a batch size of 64. All code is implemented in Mathematica.
F.1 The Gaussian probability flow ODE
I follow the standard formalism [24] with implementation improvements discussed in [28]. The forward process is given as an OU process with $\mathbf{f}(\mathbf{x},t)=\frac{1}{2}\mathbf{x}$ and $g(t)=1$ in Eq. (1) and $\boldsymbol{\xi}$ Gaussian white noise. Rescaling the standard score function in Eq. (4) then yields the loss function
$J_{\rm G}(\boldsymbol{\theta})=\mathbb{E}_{t\sim{\rm U}(0,T)}\mathbb{E}_{\mathbf{Y}_{0}\sim p_{\rm data}}\mathbb{E}_{\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\left\|\hat{\mathbf{s}}_{\boldsymbol{\theta}}\left(\mathbf{Y}_{0}e^{-t/2}+\sqrt{D(1-e^{-t})}\,\boldsymbol{\epsilon},t\right)-\boldsymbol{\epsilon}\right\|^{2}\right]$ (96)
implying a weight function $w(t)=\left(1-e^{-t}\right)/T$. The generative process is given by the probability flow ODE in exponential integrator form [28]
$\mathbf{X}_{i+1}=\mathbf{X}_{i}\,e^{\Delta t/2}-\left(e^{\Delta t/2}\sqrt{1-e^{-(T-t)}}-\sqrt{1-e^{\Delta t-(T-t)}}\right)\sqrt{D}\,\hat{\mathbf{s}}_{\boldsymbol{\theta}}(\mathbf{X}_{i},T-t)$ (97)
and the starting point is distributed as $\mathbf{X}_{0}\sim\mathcal{N}(\mathbf{0},D\mathbf{I})$.
F.2 The LIM SDE
As a variant of LIM [27] I consider the forward process as a generalized OU process driven by isotropic $\alpha$-stable noise with $1<\alpha\leq 2$
$\dot{\mathbf{Y}}(t)=-\frac{1}{\alpha}\,\mathbf{Y}(t)+\left(\frac{\sigma^{2}}{2}\right)^{1/\alpha}\boldsymbol{\xi}_{\alpha}(t)$ (98)
where the noise increments $\Delta\boldsymbol{\xi}_{\alpha}\sim S\alpha S(\Delta t)$ and the isotropic $\alpha$-stable distribution $S\alpha S(\nu)$ is described by the characteristic function
$\chi_{\alpha}(\mathbf{k})=e^{-\nu\,k^{\alpha}}$ (99)
The factor $\left(\frac{\sigma^{2}}{2}\right)^{1/\alpha}$ is introduced as a scale parameter such that in the Gaussian limit $\alpha=2$ the noise intensity is given by $\sigma^{2}$. As shown in [27] a loss function for the score function estimation is given by
$J_{\rm LIM}(\boldsymbol{\theta})=\mathbb{E}_{t\sim{\rm U}(0,T)}\mathbb{E}_{\mathbf{Y}_{0}\sim p_{\rm data}}\mathbb{E}_{\boldsymbol{\epsilon}\sim S\alpha S(1)}\left[\left\|\hat{\mathbf{s}}_{\boldsymbol{\theta}}\left(\mathbf{Y}_{0}e^{-t/\alpha}+\gamma(t)\,\boldsymbol{\epsilon},t\right)-\boldsymbol{\epsilon}\right\|^{2}\right]$ (100)
where $\gamma(t)=\left(\frac{\sigma^{2}}{2}\left(1-e^{-t}\right)\right)^{1/\alpha}$. The SDE of the generative process follows in the exponential integrator form
$\mathbf{X}_{i+1}=\mathbf{X}_{i}\,e^{\Delta t/\alpha}-\alpha\,\frac{\sigma^{2}\left(e^{\Delta t/\alpha}-1\right)}{2\gamma(T-t)^{\alpha-1}}\,\hat{\mathbf{s}}_{\boldsymbol{\theta}}(\mathbf{X}_{i},T-t)+\left(\frac{\sigma^{2}}{2}\left(e^{\Delta t}-1\right)\right)^{1/\alpha}\boldsymbol{\epsilon}$ (101)
where $\boldsymbol{\epsilon}\sim S\alpha S(1)$ and the starting point is distributed as $\mathbf{X}_{0}\sim S\alpha S\left(\frac{\sigma^{2}}{2}\right)$.
Appendix G Alternative derivation of the continuity equation for $\tilde{p}$
@@
$\tilde{\mathcal{L}}u(\mathbf{x})$ $=$ $-\mathbf{f}(\mathbf{x},T-t){\cdot}\nabla u(\mathbf{x})$ (102)
@@
$\tilde{\mathcal{L}}u(\mathbf{x})$ $=$ $-\mathbf{f}(\mathbf{x},T-t){\cdot}\nabla u(\mathbf{x})+\lambda\,g(T-t)\int\mathrm{d}\mathbf{z}\,\rho(z)\int_{0}^{1}\mathrm{d}\gamma\,\left[\mathbf{z}{\cdot}\nabla u(\mathbf{x}+\gamma g(T-t)\mathbf{z})\right]\times$ (103)
@@
$\tilde{\mathcal{L}}^{\dagger}\tilde{p}(\mathbf{x},t)$ $=$ $-\nabla{\cdot}\left(-\mathbf{f}(\mathbf{x},T-t)+\lambda\,g(T-t)\int_{0}^{1}\mathrm{d}\gamma\int\mathrm{d}\mathbf{z}\,\mathbf{z}\,\rho(z)\right$$\times$ (105)